please solve the following questions. 1. Max Leonard, vice president of Marketing for Dysk Computer, Inc must decide whether to introduce a mid-priced version of the firms DC 6900 personal computer product line the DC 6900-X. The DC6900-X would sell for $4,000, with unit variable costs of $2,250. Projections made by an independent marketing research firm indicate that the DC6900-X would achieve a sales volume of 50,000 units next year, in its first year of commercialization. One-half of the first years volume would come from competitors personal computers and market growth. However, a consumer research study indicates that 30 percent of the DC6900-X sales volume would come from the higher-priced DC6900-0mega personal computer, which sells for $3,000 (with unit variable costs of $1,800). Another 20 percent of the DC6900-X sales volume would come from the economy-priced DC6900-Alpha personal computer, priced at $2,200 (with unit variable costs of$ 1,200). The DC6900-0mega unit volume is expected to be 40,000 units next year, and the DC6900-Alpha is expected to achieve a 60,000-unit sales level. The fixed costs of launching the DC6900-X have been forecast to be $1 million during the first year of commercialization. Should Mr. Leonard add the DC6900-X model to the line of personal computers? Why?
Answer #1

Fixed Cost of Launching DC6900-X = 1000000 $ or 1 million $
Selling Price of DC6900-X = 4000 $
Unit Variable Cost of DC6900-X = 2250$
Estimation of sales by a Marketing Research Firm = 50000 Units
Total Revenue estimated in the first year of Commercialization = 50000*(4000) =200000000= 200 Million $
Total Cost Estimated in the First Year for DC6900-X = (2250*50000 +1000000) = 113.5 Million $
Total Profit Estimates = (200 - 113.5) Million $ = 86.5 Million $

Selling Price of DC6900-Omega = 3000$
Unit Variable Cost of DC6900- Omega = 1800 $
Sales Estimate for the next Year = 40000
Profit Estimate = (3000-1800)*40000 = 48 Million $

Selling Price of DC6900-Alpha = 2200 $
Unit Variable Cost of DC6900-Alpha = 1200 $
Sales Estimate for the next Year = 60000
Profit Estimate = (2200-1200)*60000 = 60 Million $

Yes Mr. Leonard should add DC6900-X to the line of Computers, Because according to the research estimate, in the first year of commercialization , the new model will give the profit of 86.5 Million $ more than the existing models.